OpenStudy (anonymous):

1)Use the clues to find the number. Clues: The number has 3 digits and they are all different the sum of the numbers is 15 All the digits are odd and less than 9 the smallest digit is in the hundreds place and the largest digit is in the ones place 2)Michael counted 46 legs on the birds and lions in Animals Country. The number of birds was 2 more than the number of lions. How many birds did Michael see in Animal Country? How many lions did he see? (hint: birds have 2 legs and a lion has 4 legs)

OpenStudy (anonymous):

I think the digits will be 7, 5, 3 7 + 5 + 3 = 15 They all are odds.. 357 is the answer for first..

OpenStudy (anonymous):

Birds = 2 + Lions So, 2*Birds + 4*Lions = 46 2(2 + Lions) + 4*Lions = 46 4 + 6*Lions = 46 Lions = 7 Birds = 9